OpenStudy (anonymous):

a school district needs 3 teachers for every 70 children. They expect to have 14,700 children in the district next year, at this time they have 612 teachers. How many more teachers are needed?

OpenStudy (atlas):

You can do this question in 2 ways

OpenStudy (atlas):

One way is a bit tedious.........keep on subtraction 70 from 14,700 students till you get 0. For everytime you subtract add 3 for teachers (because you need 3 teachers for 70 children)

OpenStudy (atlas):

*subtacting

OpenStudy (atlas):

There is another very easy method...but first tell me do you understand the first approach

OpenStudy (anonymous):

18 more teachers are needed

OpenStudy (anonymous):

u have to find how many 70 student froups are presnet.for dis divide 14700 by 70 which comes to210 gropus. as each gropu needs 3 teachers,210 groups will need 210*3 teachers=630. now,they say dat they have 612 teachers and how many more they need...so required no. of teachers=630-612 =18teachers